OpenStudy (dj3strella):

Help me please 2. -6a + 3 = -3 ( 2a - 1 ) A. a=5 b. a= 10 c. no solution d. identity

OpenStudy (dj3strella):

@Squirrels

OpenStudy (squirrels):

@abb0t

OpenStudy (anonymous):

d

OpenStudy (dj3strella):

@CausticSyndicalist

OpenStudy (causticsyndicalist):

(2a - 1) a = 1/2 or 0.5 2 times 0.5 - 1 = 0 -6a + 3 = -3 a = 1 so possibly D.

OpenStudy (dj3strella):

thank you

OpenStudy (causticsyndicalist):

you're welcome

OpenStudy (dj3strella):

@keishayay here the question u asked me and heres ur answers